# LVMark: Robust Watermark for latent video diffusion models

### v2

@@ 4.4 Training Objectives
$\mathcal{L}_{lpips}=\text{VGG}(\mathcal{D}_{mod}({\mathcal{M}}(m),{\mathcal{E}}(v)),v),$ (4)
@@
$\mathcal{L}=$ $\lambda_{message}\mathcal{L}_{message}$ (6)
$+\lambda_{lpips}\mathcal{L}_{lpips}+\lambda_{patch}\mathcal{L}_{patch},$
where $\lambda_{message}$, $\lambda_{lpips}$, and $\lambda_{patch}$ are hyperparameters that balance three loss functions.
Figure 3: Qualitative results with baselines. We show the visual quality of generated videos with baseline watermarking methods. The first row shows the crop of each image, while the second and third rows show the video frame itself and the difference map($\times$10) between the original and each method.
Open Sora DynamiCrafter
Bit Method Bits Acc.(%) $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ tLP $\downarrow$ FVD $\downarrow$ Bits Acc.(%) $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ tLP $\downarrow$ FVD $\downarrow$
32 HiDDeN [55] 90.83 30.20 0.901 0.265 5.493 2707.9 90.48 30.19 0.887 0.187 5.410 1637.7
Blind [25] 92.05 28.88 0.888 0.251 5.615 1278.9 87.56 26.44 0.852 0.239 5.348 1390.2
Stable Signature [12] 86.32 21.71 0.681 0.285 3.307 494.8 87.65 21.39 0.541 0.162 0.511 249.2
WOUAF [18] 87.28 23.62 0.838 0.166 0.856 193.7 90.62 22.38 0.839 0.125 0.763 469.4
Ours 94.44 31.15 0.919 0.105 0.219 65.73 96.00 31.50 0.892 0.075 0.173 85.49
48 HiDDeN [55] 85.62 29.85 0.888 0.269 6.802 2731.9 86.96 29.83 0.876 0.188 5.843 1716.5
Blind [25] 87.41 27.79 0.872 0.204 4.845 1365.8 86.14 25.23 0.848 0.259 5.480 1480.1
Stable Signature [12] 78.99 20.57 0.614 0.297 4.019 688.5 81.48 20.82 0.530 0.175 0.603 262.8
WOUAF [18] 83.05 23.47 0.828 0.168 1.053 196.2 89.63 22.23 0.811 0.130 0.789 523.2
Ours 90.42 30.32 0.910 0.117 0.287 117.5 93.67 30.78 0.889 0.100 0.240 155.0
Table 1: Quantitative results. We present the quantitative results for Open-Sora [54] and DynamiCrafter [44] with various watermarking methods: HiDDeN [55], Blind [25], Stable Signature [12], WOUAF [18] and LVMark, applied. The evaluation includes image metrics: PSNR, SSIM, LPIPS, video metrics: tLP [9], FVD [38], and bit accuracy on 32 and 48-bit messages.
@@ 5.1 Dataset
Video dataset. We train our method on Panda-70M dataset [7], a large-scale collection of video frames recognized for its diversity across various content types. We randomly download 10K videos from this dataset. Following the approach of Stable Video Diffusion [3], we sample 8 frames from each video, with each frame resized to a resolution of $256\times 256$ for training.
@@
Bit Accuracy(%) $\uparrow$
None
@@
Open Sora HiDDeN [55] 90.83 90.71 62.26 75.06 71.18 90.83 90.77 90.88 61.70 61.72 -
Blind [25] 92.05 87.91 86.30 68.56 62.92 86.63 55.75 89.50 84.70 82.97 -
Stable Signature [12] 86.32 78.42 54.26 84.15 76.65 86.32 77.43 76.72 81.31 51.82 77.58
WOUAF [18] 87.28 79.79 78.62 86.94 85.66 87.28 81.91 82.19 81.24 76.21 75.33
Ours 94.44 91.85 87.10 93.21 87.94 94.44 93.88 94.40 90.21 83.97 89.69
Dynami Crafter HiDDeN [55] 90.48 90.45 50.32 78.01 82.24 90.48 88.47 90.44 62.15 61.97 -
Blind [25] 87.56 83.74 86.26 67.29 59.53 84.38 57.88 84.57 83.45 85.49 -
Stable Signature [12] 87.65 84.15 50.39 82.20 77.44 87.65 83.97 85.97 83.94 51.62 80.98
WOUAF [18] 90.62 90.59 89.03 88.41 87.99 90.62 88.61 90.50 89.70 88.42 78.36
Ours 96.00 93.42 93.02 95.18 94.04 96.00 95.48 94.95 95.05 91.91 90.23
@@
Table 2: Robustness comparisons. We present bit accuracy results of watermarking methods under various attacks, including image distortions, video distortions, combined distortions, and model distortions.
5.2 Implementation Details
Video diffusion models are generally categorized into two types: DiT [28] and U-Net [33] architectures. To evaluate our method on DiT architecture, we employ Open-Sora [54], an open-source text-to-video-generative model. For the U-Net architecture, we utilize DynamiCrafter [44], which is the notable superior in video quality for text-image-to-video. We follow the same approach as DynamiCrafter by collecting images. Notably, these images are exclusively utilized in DynamiCrafter to generate videos for our evaluation process. For training, we only utilize the latent decoder of these models. Our method is trained on 10K videos (8 frames, $256\times 256$ resolution) from Panda-70M dataset [7]. We modulate 50 % of the latent decoder layers to embed the message. For optimization, we set $\lambda_{message}=0.8$, $\lambda_{lpips}=0.7$, $\lambda_{patch}=10$ in Eq. 6. Furthermore, we employ 100 random messages.
5.3 Evaluation
[…]
Baseline. To the best of our knowledge, there is no watermarking method for the video diffusion models. Therefore, we compare with four strategies for a fair comparison: 1) HiDDeN [55], is the first deep learning image watermarking method. We embed the same message into each video frame. 2) Blind [25], is the traditional video watermarking method, using 3D Discrete Wavelet Transform. We follow the implementation in Blind [25]. 3) Stable Signature [12], is the watermarking method for the text-to-image diffusion model. 4) WOUAF [18], is the state-of-the-art watermarking method for the text-to-image diffusion model. For both Stable Signature and WOUAF, we train Open-Sora [54] and DynamiCrafter [44] using the same training set as our method. Additionally, we decode the message from each video frame for HiDDeN, Stable Signature, and WOUAF.
5.4 Experimental results
Video Quality and Bit Accuracy. There is a trade-off between video quality and bit accuracy, making it a key challenge in digital watermarking to improve both simultaneously. We first compare the generated video quality and bit accuracy of our method against other approaches, as shown in Figure 3 and Table 1. We observe that Stable Signature [12] and WOUAF [18] fail to generate high-quality videos with an invisible message when extended to video generation. HiDDeN [55] does not fully address temporal consistency, as it independently embeds the invisible message into each frame of the generated video. Though Blind [25] is a traditional video watermarking method, it fails to ensure temporal consistency on synthetic videos, as demonstrated by tLP and FVD evaluations, highlighting its limitations in handling generated videos. In contrast, as shown in Figure 3, and Table 1, our method preserves video quality, including temporal consistency, while achieving the highest bit accuracy. To illustrate the differences between the original and watermarked video, we present the normalized pixel intensity difference in Figure 4, following previous watermarking visualization methods [5, 30, 13]. Our method shows the smallest difference, demonstrating superior visual quality (more details in Section 5.5).
Figure 4: Visualization of pixel intensity difference. We show the normalized pixel intensity difference between original and watermarked videos. The red line depicts our results. A lower value on the y-axis indicates greater similarity to the original video.
Video Quality and Capacity. In this section, we analyze video quality and bit accuracy based on the length of the embedded message. We compare our method with others for 32-bit and 48-bit messages, using 100 different messages. As shown in Table 1, all methods exhibit a decrease in bit accuracy as the message length increases. We note that our method achieves the highest bit accuracy compared to others, with over 90% bit accuracy at 48 bits. Furthermore, while all methods experience a noticeable drop in video quality as message length increases, our approach consistently shows a smaller decline across all metrics than other state-of-the-art methods. Our method also allows for embedding more messages in videos of comparable quality, outperforming others in both bit accuracy and video quality.
Robustness for Video Distortion. One of the priorities in digital watermarking is robustness against various distortions. In this section, we compare the robustness of our method with other approaches to demonstrate its effectiveness. As shown in Table 2, HiDDeN [55] is vulnerable to H.264 compression, while Blind [25] exhibits weaknesses against frame averaging, Gaussian blur, and JPEG compression. Additionally, Stable Signature [12] and WOUAF [18] fail to preserve the message when videos generated by the DiT model undergo rotation. In contrast, our method preserves the message across all attack metrics, thanks to the watermark decoder using 3D wavelet features.
Notably, it demonstrates exceptional robustness against video compression, especially H.264. We note that since all methods except Blind [25] are applied independently per-frame, it is naturally robust against frame drop.
Robustness for Model Distortion. We explore the scenario in which a malicious user, recognizing the existence of embedded messages in videos generated by the latent decoder, distorts the latent decoder to remove the embedded messages. We adopt a strategy that involves adding Gaussian noise ($\sigma=0.05$) to the parameters of the latent decoder. We randomly select 20 % parameters of the latent decoder. Table 1 shows that our method ensures the presence of messages even when the latent decoder is distorted.
@@ 5.5 Ablation study
For the ablation study, we use the same configuration as described in Section 5.2, with 32-bit messages.
Figure 5: The impact of weight modulation rate. Each point represents the metrics obtained by training with a specific modulation rate. We experiment with modulation rates of 0%, 25%, 50%, 75%, and 100%.
@@
Open Sora 2D DWT 92.01 30.92 0.450 122.9
3D DWT (Ours) 94.44 31.15 0.219 65.73
Dynami Crafter 2D DWT 91.71 28.73 0.383 158.3
3D DWT (Ours) 96.00 31.50 0.173 85.49
Table 3: Comparison of different DWT types. We evaluate the metrics of 2D DWT and 3D DWT.
@@
Domain Type PSNR $\uparrow$ None H264(crf21)
Open Sora RGB only 30.64 89.85 87.83
High Freq only 31.91 50.45 50.33
Low Freq only 33.30 50.13 50.12
High Freq + RGB 28.97 88.72 86.14
Low Freq + RGB (Ours) 31.15 94.44 90.21
Dynami Crafter RGB only 28.62 91.82 88.10
High Freq only 31.16 50.38 56.25
Low Freq only 29.79 51.04 50.19
High Freq + RGB 28.07 90.10 88.95
Low Freq + RGB (Ours) 31.50 96.00 95.05
Table 4: Comparison of different domain types. We evaluate the metrics for each domain type.
Weight modulation for watermarking As mentioned in Section 4.1, we selectively modulate the weights of the latent decoder to embed messages. The modulation rate is an important factor, as there exists trade-off. As shown in Figure 5, we generate videos with $k$ ranging from 0 to 100. As $k$ increases, bit accuracy improves while video quality declines, illustrating a trade-off. We find an optimal balance between bit accuracy and video quality when $k$ is set to 50.
Watermark Decoder with 3D DWT. Temporal consistency is a crucial factor in video watermarking and generation tasks. To address this, we employ the 3D Wavelet Transform (3D DWT), which captures spatio-temporal information and converts videos into wavelet signals. Table 3 shows the comparison between 3D DWT and 2D Discrete Wavelet Transform (2D DWT), which utilizes only spatial information. As shown in Table 3, our watermark decoder utilizing 3D DWT achieves the best temporal consistency, bit accuracy, and video quality.
Watermark Decoder with Low Frequency. As mentioned in Section 4.3, we initially utilize only the low frequency subband ($LLL$) among the 3D wavelet subbands. While this achieves high PSNR values, relying solely on the frequency domain limits bit accuracy, as shown in Table 4. By combining RGB features with the low frequency subband ($LLL$), we significantly enhance bit accuracy without compromising PSNR. This approach is particularly robust for decoding messages, even in H.264 compressed videos.
Figure 6: Visual impact of weighted patch loss. We visualize a video frame trained with and without weighted patch loss. To enhance clarity, we crop the local regions containing artifacts.
@@
Bits Acc. $\uparrow$ LPIPS $\downarrow$ Bits Acc. $\uparrow$ LPIPS $\downarrow$
W/o Patch Loss 95.96 0.151 96.41 0.115
W/ Patch Loss (Ours) 94.44 0.105 96.00 0.075
Table 5: Ablation on weighted patch loss. We evaluate the impact of weighted patch loss.
Impact of Weighted Patch Loss on the video. As depicted in Section. 4.4, artifacts appear in localized areas of the generated videos. To address this issue, we design a loss function that focuses on the regions where artifacts occur. Without the patch loss, although we achieve higher bit accuracy, the visual quality significantly degrades due to the presence of local artifacts, especially in the DiT model: Open-sora, which relatively focuses less on the local area by the nature of transformer architectures. Table 5 and Figure 6 demonstrate the effectiveness of the weighted patch loss in eliminating artifacts from local regions.

### v3

@@ 3.4 Training Objectives
$\mathcal{L}_{vgg}=\text{VGG}(\mathcal{D}_{mod}({\mathcal{M}}(m),{\mathcal{E}}(v)),v),$ (4)
@@
$\mathcal{L}=\lambda_{msg}\mathcal{L}_{msg}+\lambda_{vgg}\mathcal{L}_{vgg}+\lambda_{patch}\mathcal{L}_{patch},$ (6)
where $\lambda_{msg}$, $\lambda_{vgg}$, and $\lambda_{patch}$ are hyper-parameters that balance three loss functions.
Figure 3: Qualitative results with baselines. We show the visual quality of generated videos with baseline watermarking methods. The first row shows the crop of each image, while the second and third rows show the video frame itself and the difference map($\times$10) between the original and each method. Note that, unlike other approaches, VideoShield [12] embeds watermarks in the latent noise of the diffusion model, generating a video different from the original.
@@ 4.1 Dataset
Training Dataset. We train our method on Panda-70M dataset [6], a large-scale collection of video frames recognized for its diversity across various content types. We randomly download 10K videos from this dataset. Following the approach of Stable Video Diffusion [3], we sample 8 frames from each video, with each frame resized to a resolution of $256\times 256$ for training.
@@
4.2 Implementation Details
Video diffusion models are generally classified into two categories: DiT-based [26] and U-Net-based [27] architectures. To evaluate our method, we utilize Open-Sora [45] for the DiT architecture and DynamiCrafter [37] for the U-Net architecture, both of which are state-of-the-art open-source models, with Open-Sora excelling in text-to-video generation and DynamiCrafter in text-and-image-to-video generation. Our method is trained for over 40K iterations using the AdamW optimizer [20] on a single A100 GPU. We modulate 50% of the latent decoder layers to embed the message. For fair comparisons, we apply Bose–Chaudhuri–Hocquenghem (BCH) codes, a type of error correction code, to all baseline watermarking methods under the same setting to detect and correct minor bit errors. For more details, please refer to the supplementary material. The generated videos have a resolution of $512\times 320$ for DynamiCrafter and $640\times 640$ for Open-Sora, with both models generating 16 frames. For optimization, we set $\lambda_{msg}=0.8$, $\lambda_{vgg}=0.7$, $\lambda_{patch}=10$ in Eq. 6. Furthermore, we employ 100 random messages.
4.3 Evaluation
[…]
Baselines. To the best of our knowledge, there is no watermarking method for video diffusion models that preserves the original content. Therefore, for a fair comparison, we evaluate seven methods across three types of watermarking methods: 1) HiDDeN [46] and WAM [30] are watermarking methods for images. 2) Blind [23] and REVMark [44] are watermarking methods for videos. 3) Stable Signature [10], WOUAF [16], and VideoShield [12] are watermarking methods for diffusion models. For both Stable Signature and WOUAF, we train Open-Sora [45] and DynamiCrafter [37] using the same training set as our method. Additionally, we decode the message from each video frame for image-level watermarking methods [46, 30, 10, 16].
4.4 Results
Video Quality and Bit Accuracy. There is a trade-off between video quality and bit accuracy, making it a key challenge in digital watermarking to improve both simultaneously. Consequently, for a fair comparison, we measure the video quality when bit accuracy is maintained close to 100%. We first compare the generated video quality and bit accuracy of our method against other approaches, as shown in Figure 3 and Table 1. We observe that image-based watermarking methods [46, 30, 10, 16] do not fully address temporal consistency, as they independently embed the invisible message into each frame of the generated video. The performance drop of image diffusion model watermarking methods [10, 16] is particularly pronounced when applied to video diffusion models, as they are designed for image-based models that ignore temporal information. Video watermarking methods [23, 44] maintain relatively high bit accuracy and video quality but exhibit inferior performance compared to our method. Our method preserves video quality, including temporal consistency, while achieving the highest bit accuracy (more details in Section 4.5).
Watermark Capacity. In this section, we analyze the video quality and bit accuracy based on the length of the embedded message. A higher watermark bit capacity enables the registration and tracking of more users’ copyright information, enhancing the practicality of watermarking methods. We compare our method with others using 100 different messages for 32-bit and 96-bit capacities. Furthermore, we embed a 512-bit message into our method, demonstrating its ability to maintain both video quality and bit accuracy at higher capacities. As shown in Table 1, all methods exhibit a decline in bit accuracy and visual quality as the message length increases. Notably, our method achieves superior video quality, even when embedding a 512-bit message, outperforming other methods with a 32-bit capacity. Therefore, experiments with higher capacities for other methods are unnecessary. This highlights the superiority of our method in embedding high-capacity watermarks while preserving video quality better than other approaches.
Open Sora DynamiCrafter
Method Capacity (Bit) Bits Acc.(%) $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ tLP $\downarrow$ FVD $\downarrow$ Bits Acc.(%) $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ tLP $\downarrow$ FVD $\downarrow$
HiDDeN [46] 32 99.78 30.19 0.877 0.237 5.310 2980.7 99.82 30.83 0.888 0.178 7.596 1947.7
WAM [30] 32 99.99 30.84 0.917 0.170 2.738 435.4 99.99 30.92 0.911 0.156 3.777 317.1
Blind [23] 32 99.71 29.88 0.888 0.181 2.615 363.2 99.63 29.45 0.882 0.179 2.348 389.5
96 99.12 28.32 0.842 0.194 2.937 401.2 99.05 28.05 0.854 0.187 3.019 416.3
REVMark [44] 32 99.99 29.26 0.905 0.128 0.317 133.6 99.99 29.02 0.913 0.113 0.548 130.9
96 99.98 29.24 0.891 0.134 0.497 249.1 99.99 28.85 0.898 0.114 0.631 189.1
Stable Signature [10] 32 99.96 27.33 0.821 0.211 3.387 1094.8 99.02 26.78 0.797 0.223 4.353 954.8
96 95.88 26.46 0.803 0.259 4.019 1688.5 95.68 26.32 0.788 0.240 4.603 1262.8
WOUAF [16] 32 99.28 28.62 0.873 0.166 4.856 993.7 99.62 28.38 0.881 0.125 5.763 1469.4
96 96.25 28.13 0.858 0.177 5.053 1196.2 96.22 28.23 0.877 0.130 5.989 1523.2
VideoShield [12] 512 99.52 - - - - - 99.40 - - - - -
LVMark (Ours) 32 99.99 32.65 0.938 0.094 0.167 63.70 99.99 32.22 0.932 0.058 0.141 43.27
96 99.96 32.60 0.935 0.099 0.198 70.46 99.99 32.06 0.925 0.063 0.169 46.83
512 99.81 31.74 0.918 0.105 0.229 83.98 99.98 31.63 0.914 0.066 0.189 68.94
Table 1: Quantitative results. We present the quantitative results for Open-Sora [45] and DynamiCrafter [37] with various watermarking methods: image watermarking methods [46, 30], video watermarking methods [23, 44], diffusion model watermarking methods [10, 16, 12]. The evaluation includes image metrics: PSNR, SSIM, LPIPS, video metrics: tLP [7], FVD [33], and bit accuracy on 32, 96 and 512-bit messages. Note that since VideoShield [12] generates a different video from the original, the visual quality metrics in the table are not meaningful and are left blank. Best performance in bold, second in underline, lower in gray.
Capacity Image Attack Video Attack Model Attack
(Bit) None
@@ 4.4 Results
Prune / (PSNR)
(p=0.1)
Retrain / (PSNR)
Open Sora HiDDeN [46] 32 99.10 99.07 50.24 63.89 59.47 99.10 98.93 99.75 55.66 55.03 - - -
WAM [30] 32 99.99 99.94 53.74 99.95 91.94 99.99 99.97 99.99 87.60 58.03 - - -
Blind [23] 96 99.12 88.51 87.32 69.24 67.53 89.51 59.53 90.12 84.13 88.19 - - -
REVMark [44] 96 99.98 82.91 59.47 99.59 94.65 99.74 99.98 98.09 93.49 81.20 - - -
Stable Signature [10] 96 95.88 92.49 94.15 89.80 83.61 95.88 93.29 94.14 82.63 76.32 93.15 / (24.67) 93.99 / (20.03) 50.19 / (26.51)
WOUAF [16] 96 96.25 88.51 95.33 87.80 87.75 96.25 94.56 93.26 90.15 86.32 93.68 / (26.45) 92.93 / (19.67) 85.47 / (28.15)
VideoShield [12] 512 99.52 96.93 52.89 99.49 98.12 99.34 99.64 93.16 97.92 75.25 96.32 97.19 -
LVMark (Ours) 32 99.99 99.96 99.95 99.96 99.84 99.99 99.98 99.99 99.84 99.83 99.52 / (29.02) 99.85 / (22.72) 99.16 / (32.69)
96 99.96 99.91 99.82 99.95 99.77 99.96 99.96 99.96 99.76 99.81 99.46 / (28.97) 99.84 / (22.70) 98.99 / (32.58)
512 99.81 99.74 99.51 99.66 99.19 99.81 99.04 99.24 98.20 99.19 99.29 / (28.88) 99.81 / (22.53) 97.67 / (31.72)
DynamiCrafter HiDDeN [46] 32 99.02 98.82 50.38 62.64 58.81 99.02 98.91 99.81 56.01 55.82 - - -
WAM [30] 32 99.99 99.78 63.11 99.98 94.34 99.99 99.98 99.99 85.27 64.20 - - -
Blind [23] 96 99.05 84.26 85.75 69.29 61.48 85.19 57.25 86.24 85.63 85.02 - - -
REVMark [44] 96 99.99 81.95 58.27 99.88 95.86 99.79 99.99 97.83 94.52 80.29 - - -
Stable Signature [10] 96 95.68 92.11 93.86 89.03 84.11 95.68 93.23 94.01 81.17 75.32 93.68 / (23.92) 94.15 / (19.94) 50.24 / (26.35)
WOUAF [16] 96 96.22 86.25 94.84 85.80 87.47 96.22 94.42 93.10 88.90 85.09 94.13 / (25.11) 94.07 / (18.96) 83.43 / (28.21)
VideoShield [12] 512 99.40 96.65 50.11 99.20 97.99 99.26 99.35 92.57 97.60 71.96 96.05 96.81 -
LVMark (Ours) 32 99.99 99.98 99.98 99.98 99.98 99.99 99.99 99.99 99.98 99.74 99.74 / (26.45) 99.90 / (21.24) 97.36 / (32.25)
96 99.99 99.98 99.98 99.98 99.95 99.99 99.97 99.99 99.94 99.50 99.72 / (26.41) 99.87 / (21.19) 97.24 / (32.02)
512 99.98 99.90 99.93 99.74 99.94 99.98 99.04 99.36 99.75 99.16 99.62 / (26.26) 99.76 / (21.03) 95.73 / (31.59)
@@
Table 2: Robustness comparisons. We present bit accuracy of watermarking methods under various attacks. Image and video methods [46, 30, 23, 44] are not applicable to model attacks, and VideoShield [12] does not fine-tune the diffusion model, leaving them blank.
Robustness for Video Distortions. A key priority in digital watermarking is robustness against distortions. In this section, we evaluate our method’s robustness to distortions in videos. Table 2 shows the bit accuracies of our method and baselines against various distortions. Leveraging spatio-temporal information in the 3D wavelet domain, our watermark decoder achieves high bit accuracy across all evaluated attacks. Notably, it demonstrates exceptional robustness against compressions such as H.264, by extracting messages from low-frequency components. We note that since all methods except video watermarking methods [23, 44] are applied independently per frame, it is naturally robust against frame swap.
Robustness for Model Distortions. To mitigate the risk of unauthorized model usage, we assess the robustness of our method against model distortions. Since strong Gaussian noise or layer pruning degrades video quality, unauthorized users are unlikely to apply significant distortions. Maintaining high bit accuracy despite quality degradation demonstrates the robustness of the watermark decoder. Therefore, we evaluate bit accuracy under compromised video quality. Additionally, we evaluate bit accuracy when an unauthorized user retrains the model using a specific watermark key. The right side of Table 2 reports bit accuracy and PSNR under model attacks. Our method ensures the presence of messages even when the diffusion model is distorted. Model distortion alters parameters, causing variations in the generated video details. By extracting watermarks from both the wavelet and RGB domains, and training with varying watermark keys and distortion layers, our decoder becomes resilient to input variations, ensuring robustness against model distortion.
@@ 4.5 Ablation Study
For the ablation study, we follow the same setup as in Section 4.2 with 512-bit messages. For additional experimental results, please refer to the supplementary materials.
Figure 4: The impact of weight modulation rate. Each point represents the metrics for different modulation rates: 0%, 25%, 50%, 75%, and 100%.
Importance-based Weight Modulation. As mentioned in Section 3.2, we selectively modulate the top $k$ of layers with the least impact on video quality to embed messages while minimizing visual degradation. As shown in Figure 4, as the modulation rate $k$ approaches 100%, bit accuracy improves at the cost of video quality, whereas setting $k$ as 0% results in failed bit extraction, illustrating a trade-off. We find an optimal balance between bit accuracy and video quality when $k$ is set to 50.
@@
Open Sora 3D DCT 98.74 30.73 0.249 104.8
3D FFT 97.98 31.28 0.232 85.37
2D DWT 96.89 30.97 0.410 175.4
3D DWT (Ours) 99.81 31.74 0.229 83.98
DynamiCrafter 3D DCT 99.21 30.72 0.293 114.7
3D FFT 95.86 31.42 0.197 72.35
2D DWT 96.63 30.89 0.324 143.9
3D DWT (Ours) 99.98 31.63 0.189 68.94
Table 3: Comparison of different frequency types. Evaluation of bit accuracy, PSNR, tLP, and FVD across various frequency transformation techniques.
@@
Domain Type PSNR $\uparrow$ None H264 (crf24)
Open Sora RGB only 30.87 94.87 86.99
High Freq only 33.74 50.62 50.54
Low Freq only 33.85 50.23 50.13
High Freq + RGB 30.50 95.46 89.40
Low Freq + RGB (Ours) 31.74 99.81 98.20
DynamiCrafter RGB only 30.49 94.31 85.90
High Freq only 33.19 51.29 51.21
Low Freq only 33.23 50.93 51.78
High Freq + RGB 30.35 94.78 88.93
Low Freq + RGB (Ours) 31.63 99.98 99.75
Table 4: Comparison of different domain types. Evaluation of PSNR and bit accuracy using different frequency components.
Watermark Decoder with 3D DWT. Temporal consistency is crucial in video watermarking and generation. To enhance this, we use the 3D Wavelet Transform (3D DWT), which captures spatio-temporal features by decomposing videos into wavelet representations. Table 3 compares 3D Discrete Cosine Transform, 3D Fast Fourier Transform, 3D DWT, and 2D DWT. Our experiments show that 3D DWT outperforms other methods in bit accuracy and video quality due to its superior localization in time and frequency. In contrast, 2D DWT, which only captures spatial information, results in a loss of temporal consistency, degrading video quality metrics such as tLP and FVD. Based on these findings, we utilize 3D DWT in our watermark decoder.
Watermark Decoder with Low-frequency. As mentioned in Section 3.1, we utilize only the low-frequency subband ($LLL$) among the 3D wavelet subbands. While this achieves high PSNR, relying solely on the frequency domain limits bit accuracy, as shown in Table 4. By combining RGB features with the low-frequency subband ($LLL$), we significantly enhance bit accuracy without compromising PSNR. Our method is particularly robust against compression techniques like H.264, which preserve low-frequencies.
Figure 5: Visual impact of weighted patch loss. We visualize a video frame trained with and without weighted patch loss. To enhance clarity, we crop the local regions containing artifacts. Both results used VGG Loss [8] with the default setting.
Impact of Weighted Patch Loss. As described in Section 3.4, weighted patch loss mitigates localized artifacts, enhancing overall video quality. VGG Loss [8] is often used to improve the perceptual quality of images, but it has limitations in enhancing PSNR, which measures pixel-wise differences. To address this, combining VGG Loss with Mean Absolute Error (MAE) Loss, which focuses on pixel-wise differences, can be effective. However, as shown in Figure 5, our experiments reveal that using both losses together introduces localized artifacts, as the embedded watermark in the video creates artifacts that degrade visual quality. Table 5 demonstrates that weighted patch loss enables higher bit accuracy while preserving video quality, outperforming MAE Loss. This highlights the effectiveness of weighted Patch Loss, considering the trade-off between bit accuracy and visual quality.
@@
Bits Acc. $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ Bits Acc. $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$
VGG Loss 99.78 30.86 0.107 99.86 30.74 0.071
VGG + MAE Loss 99.80 31.12 0.139 99.82 31.03 0.107
VGG + Patch Loss (Ours) 99.81 31.74 0.105 99.98 31.63 0.066
Table 5: Ablation on weighted patch loss. Evaluation of bit accuracy and visual quality across loss functions.